# LLM-Powered GUI Agents in Phone Automation: Surveying Progress and Prospects

### v1

@@ 1 Introduction
AppAgent [48], AppAgent v2 [64], Auto-GUI [65], ScreenAI [66], CoCo-agent [67], leaf, text width=34.8em ] ] [ Multi-Agent
@@
PromptRPA [71], Ask-before-Plan [72], leaf, text width=26.4em ] ] [ Scenario-Based
@@
4.1.2) [ e.g. Mobile-Agent [51], OmniParser [56], VisionDroid  [54], AppAgent [48],
MM-Navigator [60], MobileExperts [55], VisionTasker [79], AppAgent v2 [64],
GUI Narrator [62], Mobile-Agent-v2 [52], Mobile-Agent-E [70], leaf, text width=26.4em ] ] ] [ Training-based Methods
@@
MobileFlow [80], ShowUI [81], Aguvis [82], UI-TARS [83], leaf, text width=20em ] ] [ Phone UI-Specific [ UI Grounding [ e.g. MUG [84], LVG [85], UI-Hawk [86],
Aria-UI [87], OS-Atlas [88], GUI-Bee [89], leaf, text width=13.7em ] ] [ UI Referring [ e.g. Ferret-UI [90], UI-Hawk [86],
@@
Answering [ e.g. ScreenAI [66], WebVLN [93], UI-Hawk [86], leaf, text width=13.7em ] ] ] ] [ Supervised Fine-Tuning
4.2.2) [ e.g. SeeClick [94], InfiGUIAgent [95], GUICourse [96], Agent-R [97], GUI Odyssey [98],
TinyClick [99], MobileAgent [100], ReALM [101], IconDesc [102], leaf, text width=26.4em ] ] [ Reinforcement Learning
4.2.3) [ Phone Agents [ e.g. DigiRL [103], DistRL [104], AutoGLM [105], ReachAgent [106], leaf, text width=20em ] ] [ Web Agents [ e.g. ETO [107], Agent Q [108], AutoWebGLM [109], GLAINTEL [110], leaf, text width=20em ] ] [ PC Agents [ e.g. ScreenAgent [111], AssistGUI [112], leaf, text width=20em ] ] ] ] ] [ Datasets and
@@
AUITestAgent [131], AgentStudio [132], AndroidLab [133], MobileAgentBench [134], VisualAgentBench [135], A3 [136], leaf, text width=34.8em ] ] ] ] ]
@@ 3.2.1 Storage
  • Knowledge Injection. Agents can enhance their decision-making by incorporating knowledge derived from exploratory interactions and stored contextual information. This involves utilizing data collected during offline exploration phases or from observed human demonstrations to inform the LLM’s reasoning process. For instance, AutoDroid [50] explores app functionalities and records UI transitions in a UI Transition Graph (UTG) memory, which are then used to generate task-specific prompts for the LLM. Similarly, AppAgent [48] compiles knowledge from autonomous interactions and human demonstrations into structured documents, enabling the LLM to make informed decisions based on comprehensive UI state information and task requirements. AppAgent v2 [64] introduces a more efficient mechanism for knowledge base construction and updating. It leverages Retrieval-Augmented Generation (RAG) technology to achieve real-time dynamic updates of knowledge base information. This significantly enhances the agent’s adaptability in new environments. Other similar works include AdaptAgent [205] and others.
@@ 3.4.1 Role-Coordinated Multi-Agent
In general computer automation, Cradle [69] leverages foundational agents with general roles to achieve versatile computer control. Agents specialize in functions like command generation or state monitoring, enabling Cradle to tackle general-purpose tasks across multiple software environment. Additionally, studies such as Ask-before-Plan [72], PromptRPA [71], LUMOS [240], and WebPilot [241] also utilize general-purpose role agents to execute tasks and excel in complex tasks like planning. Among these, LUMOS provides high-quality training data and methods for future intelligent agent research.
@@ 4.2.1 Task-Specific LLM-based Agents
General-Purpose Models. The general-purpose GUI-specific LLMs are designed to handle a wide range of tasks across different applications and interfaces. They focus on enhancing direct GUI interaction, high-resolution visual recognition, and comprehensive perception to improve the capabilities of AI agents in understanding and navigating complex mobile GUIs. One significant challenge in this domain is enabling agents to interact directly with GUIs without relying on environment parsing or application-specific APIs, which can introduce inefficiencies and error propagation. To tackle this, Auto-GUI [65] presents a multimodal agent that directly engages with the interface. It introduces a chain-of-action technique that leverages previous action histories and future action plans, enhancing the agent’s decision-making process and leading to improved performance in GUI control tasks. High-resolution input is essential for recognizing tiny UI elements and text prevalent in GUIs. CogAgent [46] addresses this by employing both low-resolution and high-resolution image encoders within its architecture. Supporting input resolutions up to $1120\times 1120$, CogAgent effectively recognizes small page elements and text. Understanding UIs and infographics requires models to interpret complex visual languages and design principles. ScreenAI [66] improves upon existing architectures by introducing a flexible patching strategy and a novel textual representation for UIs. During pre-training, this representation teaches the model to interpret UI elements effectively. Leveraging large language models, ScreenAI automatically generates training data at scale, covering a wide spectrum of tasks in UI and infographic understanding. Enhancing both perception and action response is crucial for comprehensive GUI automation. CoCo-Agent [67] proposes two novel approaches: comprehensive environment perception (CEP) and conditional action prediction (CAP). CEP enhances GUI perception through multiple aspects, including visual channels (screenshots and detailed layouts) and textual channels (historical actions). CAP decomposes action prediction into determining the action type first, then identifying the action target conditioned on the action type. Addressing the need for effective GUI agents in applications featuring extensive Mandarin content, MobileFlow [80] introduces a multimodal LLM specifically designed for mobile GUI agents. MobileFlow employs a hybrid visual encoder trained on a vast array of GUI pages, enabling it to extract and comprehend information across diverse interfaces. The model incorporates a Mixture of Experts (MoE) and specialized modality alignment training tailored for GUIs. ShowUI [81] employs the UI-Guided visual tokens selection method, which randomly selects a subset of tokens from each component during training. This approach retains the original positional information while reducing redundant tokens by 33%, thereby accelerating training speed by 1.4 times. Furthermore, by using interleaved vision-language-action streaming combined with high-quality training data, it significantly improves the training speed and performance of GUI visual agents. Aguvis [82] employs a two-stage training method to enhance the generalization and efficiency of GUI agents. It uses single-step task data to train the model’s grounding abilities and multi-step task data to develop the model’s planning and reasoning capabilities. This approach significantly improves the overall performance of the agents. UI-TARS [83] employs a more in-depth and structurally robust System-2 reasoning method, combined with online bootstrapping and reflection tuning strategies. This combination effectively assists the model in handling complex tasks in dynamic environments and continuously optimizes overall performance. Collectively, these general-purpose Task Specific Model Architectures address key challenges in phone automation by enhancing direct GUI interaction, high-resolution visual recognition, comprehensive environment perception, and conditional action prediction. By leveraging multimodal inputs and innovative architectural designs, these models significantly advance the capabilities of AI agents in understanding and navigating complex mobile GUIs, paving the way for more intelligent and autonomous phone automation solutions.
@@
  • Screen Question Answering involves answering questions about the content and functionality of a screen based on visual and textual information. This task requires agents to comprehend complex screen layouts and extract relevant information to provide accurate answers. ScreenAI [66] specializes in understanding screen UIs and infographics, leveraging the common visual language and design principles shared between them. By introducing a flexible patching strategy and a novel textual representation for UIs, ScreenAI pre-trains models to interpret UI elements effectively. Using large language models to automatically generate training data, ScreenAI covers tasks such as screen annotation and screen QA. WebVLN [93] extends vision-and-language navigation to websites, where agents navigate based on question-based instructions and answer questions using information extracted from target web pages. By integrating visual inputs, linguistic instructions, and web-specific content like HTML, WebVLN enables agents to understand both the visual layout and underlying structure of web pages, enhancing screen QA capabilities. UI-Hawk [86] further enhances screen QA by enabling agents to process screen sequences and answer questions based on historical interactions. By incorporating screen question answering as one of its fundamental tasks, UI-Hawk improves the agent’s ability to comprehend and reason about screen content over time.
@@ 4.2.3 Reinforcement Learning
Phone Agents. Training phone GUI agents using RL presents unique challenges due to the dynamic and complex nature of mobile applications. Agents must adapt to real-world stochasticity and handle the intricacies of interacting with diverse mobile environment. Recent works have addressed these challenges by developing RL frameworks that enable agents to learn from interactions and improve over time. DigiRL [103] andDistRL [104] both tackle the limitations of pre-trained vision-language models (VLMs) in decision-making tasks for device control through GUIs. Recognizing that static demonstrations are insufficient due to the dynamic nature of real-world mobile environment, these works introduce RL approaches to train agents capable of in-the-wild device control. DigiRL proposes an autonomous RL framework that employs a two-stage training process: an initial offline RL phase to initialize the agent using existing data, followed by an offline-to-online RL phase that fine-tunes the model based on its own interactions. By building a scalable Android learning environment with a VLM-based evaluator, DigiRL identifies key design choices for effective RL in mobile GUI domains. The agent learns to handle real-world stochasticity and dynamism, achieving significant improvements over supervised fine-tuning, with a 49.5% absolute increase in success rate on the Android-in-the-Wild dataset. Similarly, DistRL introduces an asynchronous distributed RL framework specifically designed for on-device control agents on mobile devices. To address inefficiencies in online fine-tuning and the challenges posed by dynamic mobile environment, DistRL employs centralized training and decentralized data acquisition. Leveraging an off-policy RL algorithm tailored for distributed and asynchronous data utilization, DistRL improves training efficiency and agent performance by prioritizing significant experiences and encouraging exploration. Experiments show that DistRL achieves a 20% relative improvement in success rate compared to state-of-the-art methods on general Android tasks. Building upon these advancements, AutoGLM [105] extends the application of RL to both phone and web platforms. AutoGLM presents a series of foundation agents based on the ChatGLM model family, aiming to serve as autonomous agents for GUI control. A key insight from this work is the design of an intermediate interface that separates planning and grounding behaviors, allowing for more agile development and enhanced performance. By employing self-evolving online curriculum RL, AutoGLM enables agents to learn from environmental interactions and adapt to dynamic GUI environment. The approach demonstrates impressive success rates on various benchmarks, showcasing the potential of RL in creating versatile GUI agents across platforms. ReachAgent [106] decomposes mobile agent tasks into two sub-tasks: page reaching and page operation, utilizing a two-stage fine-tuning strategy. In the first stage, supervised fine-tuning enables the agent to better perform each sub-task. In the second stage, reinforcement learning is applied to further optimize the agent’s overall task completion capabilities, thereby enhancing its performance in complex tasks.

### v2

@@ 1 Introduction
AppAgent [48], AppAgent v2 [64], AppAgentX [65], Auto-GUI [66], ScreenAI [67], Mobile-Agent-v [68],
OS-Kairos [69], GUI-Xplore [70], CoCo-agent [71], leaf, text width=34.8em ] ] [ Multi-Agent
@@
PromptRPA [75], CHOP [76], Agent S2 [77], Ask-before-Plan [78], leaf, text width=26.4em ] ] [ Scenario-Based
@@
4.1.2) [ e.g. Mobile-Agent [51], Mobile-Agent-v2 [52], OmniParser [56], VisionDroid  [54],
AppAgent [48], MM-Navigator [60], MobileExperts [55], VisionTasker [85],
AppAgent v2 [64], GUI Narrator [62], ReuseDroid [86], VLM-Fuzzer [87],
Mobile-Agent-E [74], leaf, text width=26.4em ] ] ] [ Training-based Methods
@@
MobileFlow [88], ShowUI [89], Aguvis [90], ViMo [91], UI-TARS [92], leaf, text width=20em ] ] [ Phone UI-Specific [ UI Grounding [ e.g. MUG [93], LVG [94], UI-Hawk [95],
Aria-UI [96], OS-Atlas [97], MP-GUI [98],
Smoothing Grounding [99], GUI-Bee [100], leaf, text width=13.7em ] ] [ UI Referring [ e.g. Ferret-UI [101], UI-Hawk [95], MP-GUI [98],
@@
Answering [ e.g. ScreenAI [67], WebVLN [104], MP-GUI [98],
UI-Hawk [95], leaf, text width=13.7em ] ] ] ] [ Supervised Fine-Tuning
4.2.2) [ e.g. SeeClick [105], InfiGUIAgent [106], GUICourse [107], Agent-R [108],
GUI Odyssey [109], TinyClick [110], MobileAgent [111], ReALM [112], AppVLM [113],
V-Droid [114], IconDesc [115], leaf, text width=26.4em ] ] [ Reinforcement Learning
4.2.3) [ Phone Agents [ e.g. DigiRL [116], DistRL [117], AutoGLM [118], Digi-q [119],
ReachAgent [120], VSC-RL [121], Ui-r1 [122],
GUI-R1 [123], leaf, text width=20em ] ] [ Web Agents [ e.g. ETO [124], Agent Q [125], AutoWebGLM [126], GLAINTEL [127], leaf, text width=20em ] ] [ PC Agents [ e.g. ScreenAgent [128], AssistGUI [129], leaf, text width=20em ] ] ] ] ] [ Datasets and
@@
AUITestAgent [148], AgentStudio [149], AndroidLab [150], MobileAgentBench [151], VisualAgentBench [152],
FedMABench [153], AutoEval [154], LearnAct [155], A3 [156], leaf, text width=34.8em ] ] ] ] ]
@@ 3.2.1 Storage
  • Knowledge Injection. Agents can enhance their decision-making by incorporating knowledge derived from exploratory interactions and stored contextual information. This involves utilizing data collected during offline exploration phases or from observed human demonstrations to inform the LLM’s reasoning process. For instance, AutoDroid [50] explores app functionalities and records UI transitions in a UI Transition Graph (UTG) memory, which are then used to generate task-specific prompts for the LLM. Similarly, AppAgent [48] compiles knowledge from autonomous interactions and human demonstrations into structured documents, enabling the LLM to make informed decisions based on comprehensive UI state information and task requirements. AppAgent v2 [64] introduces a more efficient mechanism for knowledge base construction and updating. It leverages Retrieval-Augmented Generation (RAG) technology to achieve real-time dynamic updates of knowledge base information. This significantly enhances the agent’s adaptability in new environments.AppAgentX [65] introduces an evolutionary mechanism that enables dynamic learning from past interactions and replaces inefficient low-level operations with high-level actions. Other similar works include AdaptAgent [225], Mobile-Agent-V [68], LearnAct [155] and others.
@@ 3.4.1 Role-Coordinated Multi-Agent
CHOP [76] introduces a mobile operating assistant with Constrained High-frequency Optimized subtask Planning. This approach addresses challenges in the subtask level, which links high-level goals with low-level executable actions. CHOP overcomes VLM’s deficiency in GUI scenario planning by using human-planned subtasks as basis vectors, significantly improving both effectiveness and efficiency across multiple applications in both English and Chinese contexts. The framework specifically targets two common issues: ineffective subtasks that lower-level agents cannot execute and inefficient subtasks that fail to contribute to higher-level task completion.
In general computer automation, Cradle [73] leverages foundational agents with general roles to achieve versatile computer control. Agents specialize in functions like command generation or state monitoring, enabling Cradle to tackle general-purpose tasks across multiple software environment. Additionally, studies such as Ask-before-Plan [78], PromptRPA [75], LUMOS [260], and WebPilot [261] also utilize general-purpose role agents to execute tasks and excel in complex tasks like planning. Among these, LUMOS provides high-quality training data and methods for future intelligent agent research. Agent S2 [77] presents a compositional generalist-specialist framework for computer use agents that delegates cognitive responsibilities across various models. It introduces a Mixture-of-Grounding technique for precise GUI localization and Proactive Hierarchical Planning that dynamically refines action plans at multiple temporal scales based on evolving observations.
@@ 4.1 Prompt Engineering
Mobile-Agent-V [68] [Uncaptioned image] 2025.02 General GPT-4o Coordinate
LearnAct [155] [Uncaptioned image] 2025.02 General Gimini-1.5 Coordinate
@@ 4.2 Training-Based Models
V-Droid [114] 2025.03 General Llama-3.1-8B 8b Verifier-driven framework
MP-GUI [98] [Uncaptioned image] 2025.03 General InternVL2-8B 8B Screen Under- standing, Referring
@@ 4.2.1 Task-Specific LLM-based Agents
General-Purpose Models. The general-purpose GUI-specific LLMs are designed to handle a wide range of tasks across different applications and interfaces. They focus on enhancing direct GUI interaction, high-resolution visual recognition, and comprehensive perception to improve the capabilities of AI agents in understanding and navigating complex mobile GUIs. One significant challenge in this domain is enabling agents to interact directly with GUIs without relying on environment parsing or application-specific APIs, which can introduce inefficiencies and error propagation. To tackle this, Auto-GUI [66] presents a multimodal agent that directly engages with the interface. It introduces a chain-of-action technique that leverages previous action histories and future action plans, enhancing the agent’s decision-making process and leading to improved performance in GUI control tasks. High-resolution input is essential for recognizing tiny UI elements and text prevalent in GUIs. CogAgent [46] addresses this by employing both low-resolution and high-resolution image encoders within its architecture. Supporting input resolutions up to $1120\times 1120$, CogAgent effectively recognizes small page elements and text. Understanding UIs and infographics requires models to interpret complex visual languages and design principles. ScreenAI [67] improves upon existing architectures by introducing a flexible patching strategy and a novel textual representation for UIs. During pre-training, this representation teaches the model to interpret UI elements effectively. Leveraging large language models, ScreenAI automatically generates training data at scale, covering a wide spectrum of tasks in UI and infographic understanding. Enhancing both perception and action response is crucial for comprehensive GUI automation. CoCo-Agent [71] proposes two novel approaches: comprehensive environment perception (CEP) and conditional action prediction (CAP). CEP enhances GUI perception through multiple aspects, including visual channels (screenshots and detailed layouts) and textual channels (historical actions). CAP decomposes action prediction into determining the action type first, then identifying the action target conditioned on the action type. Addressing the need for effective GUI agents in applications featuring extensive Mandarin content, MobileFlow [88] introduces a multimodal LLM specifically designed for mobile GUI agents. MobileFlow employs a hybrid visual encoder trained on a vast array of GUI pages, enabling it to extract and comprehend information across diverse interfaces. The model incorporates a Mixture of Experts (MoE) and specialized modality alignment training tailored for GUIs. ShowUI [89] employs the UI-Guided visual tokens selection method, which randomly selects a subset of tokens from each component during training. This approach retains the original positional information while reducing redundant tokens by 33%, thereby accelerating training speed by 1.4 times. Furthermore, by using interleaved vision-language-action streaming combined with high-quality training data, it significantly improves the training speed and performance of GUI visual agents. Aguvis [90] employs a two-stage training method to enhance the generalization and efficiency of GUI agents. It uses single-step task data to train the model’s grounding abilities and multi-step task data to develop the model’s planning and reasoning capabilities. This approach significantly improves the overall performance of the agents. UI-TARS [92] employs a more in-depth and structurally robust System-2 reasoning method, combined with online bootstrapping and reflection tuning strategies. This combination effectively assists the model in handling complex tasks in dynamic environments and continuously optimizes overall performance. V-Droid [114] introduces a novel verifier-driven architecture where the LLM does not generate actions directly but instead scores and selects from a finite set of extracted actions, improving task success rates and significantly reducing latency. Collectively, these general-purpose Task Specific Model Architectures address key challenges in phone automation by enhancing direct GUI interaction, high-resolution visual recognition, comprehensive environment perception, and conditional action prediction. By leveraging multimodal inputs and innovative architectural designs, these models significantly advance the capabilities of AI agents in understanding and navigating complex mobile GUIs, paving the way for more intelligent and autonomous phone automation solutions.
@@
  • Screen Question Answering involves answering questions about the content and functionality of a screen based on visual and textual information. This task requires agents to comprehend complex screen layouts and extract relevant information to provide accurate answers. ScreenAI [67] specializes in understanding screen UIs and infographics, leveraging the common visual language and design principles shared between them. By introducing a flexible patching strategy and a novel textual representation for UIs, ScreenAI pre-trains models to interpret UI elements effectively. Using large language models to automatically generate training data, ScreenAI covers tasks such as screen annotation and screen QA. WebVLN [104] extends vision-and-language navigation to websites, where agents navigate based on question-based instructions and answer questions using information extracted from target web pages. By integrating visual inputs, linguistic instructions, and web-specific content like HTML, WebVLN enables agents to understand both the visual layout and underlying structure of web pages, enhancing screen QA capabilities. UI-Hawk [95] further enhances screen QA by enabling agents to process screen sequences and answer questions based on historical interactions. By incorporating screen question answering as one of its fundamental tasks, UI-Hawk improves the agent’s ability to comprehend and reason about screen content over time. MP-GUI [98] introduces a specialized MLLM for GUI understanding with three dedicated perceivers for graphical, textual, and spatial modalities. Using a fusion gate to adaptively combine these modalities and an automated data collection pipeline to address training data scarcity, MP-GUI achieves strong performance on GUI understanding tasks including screen QA despite limited training data.
@@ 4.2.3 Reinforcement Learning
ReachAgent [120] 2025.02 Phone Hybrid RL MobileVLM [63] N/A
VEM [279] [Uncaptioned image] 2025.02 Phone Environment- Free RL N/A N/A
Digi-Q [119] [Uncaptioned image] 2025.02 Phone Q-Function Based RL N/A N/A
VSC-RL [121] [Uncaptioned image] 2025.02 Phone Variational Subgoal- Conditioned RL N/A N/A
UI-R1 [122] [Uncaptioned image] 2025.03 Phone Rule-Based RL Qwen2.5-VL 3B
@@
Phone Agents. Training phone GUI agents using RL presents unique challenges due to the dynamic and complex nature of mobile applications. Agents must adapt to real-world stochasticity and handle the intricacies of interacting with diverse mobile environment. Recent works have addressed these challenges by developing RL frameworks that enable agents to learn from interactions and improve over time. DigiRL [116] andDistRL [117] both tackle the limitations of pre-trained vision-language models (VLMs) in decision-making tasks for device control through GUIs. Recognizing that static demonstrations are insufficient due to the dynamic nature of real-world mobile environment, these works introduce RL approaches to train agents capable of in-the-wild device control. DigiRL proposes an autonomous RL framework that employs a two-stage training process: an initial offline RL phase to initialize the agent using existing data, followed by an offline-to-online RL phase that fine-tunes the model based on its own interactions. By building a scalable Android learning environment with a VLM-based evaluator, DigiRL identifies key design choices for effective RL in mobile GUI domains. The agent learns to handle real-world stochasticity and dynamism, achieving significant improvements over supervised fine-tuning, with a 49.5% absolute increase in success rate on the Android-in-the-Wild dataset. Similarly, DistRL introduces an asynchronous distributed RL framework specifically designed for on-device control agents on mobile devices. To address inefficiencies in online fine-tuning and the challenges posed by dynamic mobile environment, DistRL employs centralized training and decentralized data acquisition. Leveraging an off-policy RL algorithm tailored for distributed and asynchronous data utilization, DistRL improves training efficiency and agent performance by prioritizing significant experiences and encouraging exploration. Experiments show that DistRL achieves a 20% relative improvement in success rate compared to state-of-the-art methods on general Android tasks. Building upon these advancements, AutoGLM [118] extends the application of RL to both phone and web platforms. AutoGLM presents a series of foundation agents based on the ChatGLM model family, aiming to serve as autonomous agents for GUI control. A key insight from this work is the design of an intermediate interface that separates planning and grounding behaviors, allowing for more agile development and enhanced performance. By employing self-evolving online curriculum RL, AutoGLM enables agents to learn from environmental interactions and adapt to dynamic GUI environment. The approach demonstrates impressive success rates on various benchmarks, showcasing the potential of RL in creating versatile GUI agents across platforms.
Recent advances have brought several innovative approaches to reinforcement learning for phone GUI agents. ReachAgent [120] decomposes mobile agent tasks into two sub-tasks: page reaching and page operation, utilizing a two-stage fine-tuning strategy. In the first stage, supervised fine-tuning enables the agent to better perform each sub-task. In the second stage, reinforcement learning is applied to further optimize the agent’s overall task completion capabilities, thereby enhancing its performance in complex tasks. VEM [279] introduces an environment-free RL framework that decouples value estimation from policy optimization using a pretrained Value Environment Model. Unlike traditional RL methods that require costly environment interactions, VEM predicts state-action values directly from offline data, distilling human-like priors about GUI interaction outcomes. This approach avoids compounding errors and enhances resilience to UI changes by focusing on semantic reasoning. Digi-Q [119] presents an approach to train VLM-based action-value Q-functions for device control. Instead of using on-policy RL with actual environment rollouts, Digi-Q trains the Q-function using offline temporal-difference learning on frozen, intermediate-layer features of a VLM. This approach enhances scalability and reduces computational costs compared to fine-tuning the entire VLM. The trained Q-function then uses a Best-of-N policy extraction operator to imitate the best action without requiring environment interaction. VSC-RL [121] addresses the learning inefficiencies in tackling complex sequential decision-making tasks with sparse rewards and long-horizon dependencies. By reformulating vision-language sequential tasks as a variational goal-conditioned RL problem, VSC-RL optimizes the SubGoal Evidence Lower BOund (SGC-ELBO). This approach maximizes subgoal-conditioned return via RL while minimizing the difference with the reference policy. UI-R1 [122] explores how rule-based RL can enhance reasoning capabilities of multimodal large language models for GUI action prediction. Using a small yet high-quality dataset of 136 challenging tasks, UI-R1 introduces a unified rule-based action reward enabling model optimization via Group Relative Policy Optimization (GRPO).